Reporting Issuance and Retirement of Long-Term Debt On the basis of the details of the following bonds payable and related discount accounts, indicate the tems to be reported in the Finanding Adivities section of the statement of cash flows, assuming no gain or loss on retiring the bonds ACCOUNT Bonds Payable ACCOUNT NO. Balance Date Item Debit Credit DebitCredit an.Balance 2 Retire bonds June 30 Issue bonds 750,000 600,000 1,050,000 450,000 ACCOUNT Discount on Bonds Payable ACCOUNT NO. Balance Debit Credit Debit Credit Date Item an1 Balance 33,750 12,00021,750 51,750 2,525 49,125 2 Retire bonds June30 Issue bonds Dec. 31 Amortize discount Item Retire bonds Issue bonds Amortization of discount Section of Statement of Cash Flows Added or Deducted Amount
Answer #1

Cash proceeds from the issuance of bonds = 450,000 - 30,000 = $420,000

Cash Paid to redeem bonds = 150,000 - 12,000 = $138,000

Item Section of statement of cash flows Added or Deducted Amount

Retire Bonds Financing activities Deducted $138,000

Issue Bonds   Financing activities Added   $420,000

Amortisation of Discount Operating activities Added $2,625